Please help me with the question
The distance between the bases on Max's baseball field is 118ft. Ashley made a throw from the home plate to first base to get Mark out and Salley threw direclty to third base to get sam out. How far did the ball travel in those two throws?

Answer by rothauserc(4718)

You can use the Pythagorean formula
c^2 = 118^2 + 118^2 = 27848
c = 166.88 and
both throws = 118 + 166.88 = 284.88 feet